OpenStudy (anonymous):

which is true..2.8<2 3/5 or 1.4>1 1/5 or 1.4<1 3/20 or 2.8> 2 17/20

OpenStudy (anonymous):

2.8= 2 4/5 2 3/5= 2.6 1.4= 1 2/5 1 1/5= 1.2 1 3/20= 1.15 2 17/20 =2.85

OpenStudy (anonymous):

not sure which one is true A 2.8< 2 3/5 b 1.4 > 1 1/5 c 1.4 < 1 3/20 d 2.8 > 2 17/20

OpenStudy (anonymous):

here lemme rewrite you problem

OpenStudy (anonymous):

A 2.8< 2.6 b 1.4 > 1.2 c 1.4 < 1.15 d 2.8 > 2.85 A. is 2.8 less than 2.6? B. is 1.4 greater than 1.2? C. is 1.4 less than 1.15? D. is 2.8 greater than 2.85?

OpenStudy (anonymous):

which is greater a 0.035 b 0.35 c 0.07 d 0.7

OpenStudy (anonymous):

if it isnt obvious by looking at it, multiple all of them by 100 to get rid the decimals Which is greater A. 35 B. 350 C. 70 D.700

OpenStudy (anonymous):

i meant 1000

OpenStudy (anonymous):

multiply*

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Owen scored 92 points on a 100 point test and 27 on a 30 point quiz. On which did he get a larger fraction of the total points correct? Show your work.

OpenStudy (anonymous):

so two fractions 92/100 and 27/30 simplifying these we get 23/25 and 9/10 now we multiply each fraction by some number so that the denominators are the same so (23/25)*(4/4) = 92/100 and (9/10)*(10/10)= 90/10 so which is larger 92/100 or 90/100

OpenStudy (anonymous):

92/100

OpenStudy (anonymous):

Describe a situation that can be expressed with a two-step equation. Write the equation, solve it, and check the solution, showing all your work

OpenStudy (anonymous):

so something like 2x-3=5 ?

OpenStudy (anonymous):

The cost of a ticket to an amusement park was increased by 4. The new cost is .28 a. Write a one-step equation that could be solved to find the old price. Let p represent the old price. b. Solve the equation you wrote in Part a. Show all your work. c. Explain how you know that the solution you found in Part b is the correct answer